OpenStudy (anonymous):

1)the electrical resistance for a conductor can be found by the formula R=KL/D^2.solve the formula for L

OpenStudy (anonymous):

where is the rest of it?

OpenStudy (kainui):

@gulapipry has everything to solve this problem. I'll give you a hint, suppose you want to solve for x here: \[\Large a=bx \\ \Large \text{ } \\ \Large \frac{a}{b}=x\] We just divided both sides by b to get x by itself. If be had b on the bottom, dividing x instead, we could multiply both sides by b to get x by itself too.

OpenStudy (xapproachesinfinity):

have you tried anything after what @Kainui said? we want to isolate L on the right hand side

OpenStudy (anonymous):

yes, this one is very difficult for me

OpenStudy (xapproachesinfinity):

how about starting with dividing both side by K what would we get?

OpenStudy (xapproachesinfinity):

okay let's right that down \[\frac{R}{K}=\frac{KL}{KD^2}\] what do you think we should do from here

OpenStudy (anonymous):

cross out

OpenStudy (xapproachesinfinity):

not yet first let's take care of k

OpenStudy (anonymous):

cross multiply

OpenStudy (xapproachesinfinity):

eh you just want to cross multiply lol

OpenStudy (xapproachesinfinity):

k/k is equal to what?

OpenStudy (anonymous):

I don't know

OpenStudy (xapproachesinfinity):

what a number divided by itself equal to?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

maybe factor

OpenStudy (xapproachesinfinity):

hmm how about 2/2 for example?

OpenStudy (anonymous):

1

OpenStudy (xapproachesinfinity):

good so any number divided by itself is 1 right?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

so k/k=1 now let me ask you another question what a number multiplied by 1?

OpenStudy (anonymous):

maybe 1

OpenStudy (xapproachesinfinity):

no quite any number multiplied by 1 we get that same number back

OpenStudy (anonymous):

then

OpenStudy (xapproachesinfinity):

so according to this we should get \[\frac{R}{K}=\frac{L}{D^2}\]

OpenStudy (xapproachesinfinity):

this is really not gonna make sense to you if you have trouble with arithmetic

OpenStudy (xapproachesinfinity):

so can you repeat to me what we just did

OpenStudy (anonymous):

do you really help me doing this math or you will just make fun of me. If I know that math I will not wait for 9 hours. I don't know this math so I am here for math help

OpenStudy (xapproachesinfinity):

no i'm really trying to help sorry if you felt that way, i didn't mean anything like that

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

if i'm making fun of you i wouldn't waste my time trying my hardest to explain to you

OpenStudy (xapproachesinfinity):

so did you understand what i did

OpenStudy (xapproachesinfinity):

let me tell you something homework is for learning not just to get a point and finish it what your professor is concerned about is you tried your hardest and gave it your all the things you stuck on you are supposed to bring them to your prof so he/she could explain to you.

OpenStudy (xapproachesinfinity):

when i was explaining to you what I'm really concerned about is did you understand what we are doing does it make sense to you, can you do more like this one if you get it?